# Embedding in Recommender Systems: A Survey

### v2

@@ 1.3. Survey and Tool
Tool: LibFM222libFM: https://github.com/srendle/libfm (Rendle, 2012a) is a famous toolkit to achieve FM (Rendle, 2010). LibFFM [^3] LibFFM: https://github.com/ycjuan/libffm and xlearn444xlearn: https://github.com/aksnzhy/xlearn are used to achieve FFM (Juan et al., 2016). For other algorithms, corresponding toolkits can also be found. Such as DeepCTR [^5] DeepCTR: https://github.com/shenweichen/DeepCTR, which can achieve various CTR prediction models including DeepFM (Guo et al., 2017).
1.4. Future Direction
@@
Synergistic Model Fusion for Complex Scenarios An interesting avenue for future exploration lies in the synergy between different models. Instances like CAT-E and CAT-NT, as proposed by Zhao et al. (Zhao et al., 2017), creatively combine SVD, FM, and gradient boosting, showcasing the potential of model fusion. These approaches cater to highly specialized scenarios, often tied to the designs of MF models based on SVD and FM. Investigating and expanding upon such combinations could yield even more tailored solutions for intricate use cases.
2. Self-supervised learning for embeddings
Figure 2. Data flow of SSL recommendation model
Self-supervised learning (SSL) makes use of enormous unlabeled data to train machine learning models. The preliminary idea of SSL is that use pretext tasks to collect useful and transferable information from the massive unlabeled data and leverage the pre-learned knowledge to achieve better performance on target tasks, which is formalized as follows:
(5) $z_{\theta^{*}},d_{\phi^{*}}=\underset{z_{\theta},d_{\phi}}{\arg\min\mathcal{L}_{ssl}}\left(d_{\phi}\left(z_{\theta}(\mathcal{S},\hat{\mathcal{S}})\right)\right)$
where $S$ is the original data, and $\hat{S}$ is the corrupted data (a.k.a augmented data). $z_{\theta}$ is the encoder parameter to generate the node embedding vector, which is specific for the input type. $d_{\phi}$ is the decoder to learn its own parameter $\phi$ and refine the encoder parameter $\theta$ for the main recommendation task.
Based on Eq. 5, learning embedding parameters is a part process of the self-supervised learning recommendation (SSR) model which is conducted on both pretext tasks and recommendation tasks, and the supervised signals are produced from the raw data semi-automatically. Based on the type of auxiliary tasks, we divide the SSL methods into two categories: contrastive and generative. From Fig 2, we can preliminarily see the whole picture of the SSL-based recommendation model. Furthermore, within this section, contrastive and generative methods work with either sequence data or graph data as their input. Note that this section concentrates solely on graph self-supervised learning and does not explore other graph representation learning methods. In the next section, we will offer a more comprehensive discussion of graph embedding in recommender systems.
2.1. Contrastive Learning for Embeddings
Contrastive learning (CL) is a vital research line in self-supervised learning, which aims to generate the self-supervised signal by minimizing the distance between views of the same sample and maximizing that between views of different samples in the embedding space. Different views can be generated by performing data augmentation on the training data. Corresponding to the Eq. 5, $\hat{S}$ is the set of corrupted data processed by data augmentation techniques (i.e., $\hat{S}=\mathcal{A}(S)$) which will be presented in the following paragraph.
In summary, it contains three components including 1) data augmentation (i.e., $\hat{\mathcal{S}}$); 2) sample encoder (i.e., $z_{\theta}(\cdot)$) which is particular for the type of input (e.g., graph encoder and sequence encoder); 3) CL loss (i.e., $\mathcal{L}_{CL}$). The SSR model utilizes augmentation methods on the original data and the encoder formalizes the original data and corrupted data as vectors. Finally, to discriminate the augmented data and original data, it maximizes the mutual information between the positive pairs (i.e., augmented data from the same sample) and the negative pairs (i.e., augmented data from a different sample) via the CL loss. We first introduce the augmentation techniques of graph and sequence, then demonstrate CL methods based on the type of inputs, which typically are graphs and sequences.
2.1.1. Graph Data
In the context of a graph $\mathcal{G}=(V,E)$ and adjacent matrix $\mathbf{A}$, as shown in Fig 3, three common augmentation techniques are applied as follows: Node/Edge Dropout involves removing nodes or edges with a probability $\alpha$, aiming to identify influential components for the recommendation task and enhance system robustness. Graph Diffusion introduces user-item similarity through new edges generated by weighted neighbor edge sums to model potential preferences, often selecting top-K similarities. Subgraph Sampling fabricates an augmented graph by the selection of nodes and edges, shaping a subgraph that accentuates local connectivity. These augmentation methods contribute to the self-supervision signal for learning high-quality embeddings.
SGL (Wu et al., 2021b) focuses on learning the user node and item node embedding in the user-item bipartite graph, which is the first work utilizing contrastive learning for recommendation by designing an auxiliary task to supplement the recommendation model. In particular, it first uses edge dropout, node dropout, and random walk to generate three views of the samples (e.g., user nodes or item nodes). Then, it creates positive pairs (i.e., $\left\{\left(\mathbf{A}_{s}^{\prime},\mathbf{A}_{s}^{\prime\prime}\right)\mid s\in\mathcal{S}\right\}$ ) using views of the same sample and negative pairs (i.e., $\left\{\left(\mathbf{A}_{s_{1}}^{\prime},\mathbf{A}_{s_{1}}^{\prime\prime}\right)\mid s_{1},s_{2}\in\mathcal{S},{s_{1}\neq s_{2}}\right\}$) by views of different samples. The auxiliary task ensures that different views of the same nodes are similar in the embedding space. In other words, it aims to maximize the similarity between positive pairs and minimize that of negative pairs to learn the node embeddings. Thus, following SimCLR (Chen et al., 2020a), it adopts the InfoNCE loss (Oord et al., 2018), which could maximize the agreement between positive samples and minimize the agreement between negative samples. Moreover, the node embeddings are computed by a LightGCN  (He et al., 2020), a graph neural network (GNN) acting as the embedding layer. $\tau$ is the hyper-parameter known as the temperature in the softmax function. $s\in\mathcal{S}$ is the sample node from the user nodes or the item nodes. Finally, SGL combines the user and item contrastive loss. It jointly optimizes the InfoNCE loss (i.e., for the CL task) and the Bayesian personalized ranking (BPR) loss (i.e., for the recommendation task).
HHGR (Zhang et al., 2021a) recommends items for groups in a user-item graph and tackles the issue of distorted local structures due to random node dropout. They propose a dual-scale node-dropping strategy: coarse-grained node dropout eliminates nodes across groups, while fine-grained node dropout targets nodes within the same preference group. Their approach shares similarities with the SGL method, focusing on mutual information between distinct node perspectives. In the realm of cross-domain recommendation utilizing CL, methods like CCDR (Xie et al., 2022a) transfer valuable node embeddings from a well-equipped source domain to a less-equipped target domain. CCDR employs both intra-domain and inter-domain CL tasks for knowledge transfer. In the intra-domain CL task, two neighbor sets of node $i$ are obtained, and GNN aggregation focuses on different neighbors of $i$ to obtain two embeddings ($e_{i}$, $e_{\prime(i)}$). The InfoNCE loss is then computed using the positive pair ($e_{\prime(i)},e_{i}$). The inter-domain CL task transfers knowledge across domains by conducting CL on nodes with identical properties (e.g., users and taxonomies) in both domains. Similarly, PCRec (Wang et al., 2021c) addresses cross-domain tasks in two phases. The first phase involves pre-training on the source domain to extract knowledge, and the second phase employs target domain knowledge to transfer insights to the recommendation model. DCL (Liu et al., 2021c) challenges the assumption of interest in only sampled negative items by conducting subgraph sampling through perturbation of the $L$-hop ego network of a given node.
2.1.2. Sequence Data
@@
In the context of sequence data, this section introduces four primary techniques to generate corrupted sequences to facilitate sequence augmentation. Item Cropping, inspired by image cropping in computer vision, is applied to contrastive learning in the sequential recommendation. Item Masking, borrowed from language models, involves masking elements in a sequence to prevent overfitting and encourage learning of high-order structures among tokens. Similarly, Item Reordering is proposed to ensure predictions remain invariant to item order, training recommendation models on augmented data with varying item sequences. Item Substitution enhances recommendation models by maintaining the original sequence’s structure while substituting highly correlated yet possibly redundant items. Item Insertion, akin to substitution, involves adding related items before each original item in a subset of the sequence. Specific details are shown in Fig. 4.
Figure 3. Augmentation methods of Graph
Figure 4. Augmentation methods of Sequence
@@
Contrastive methods utilize the CL loss with augmented data to enhance embeddings by capturing latent patterns. However, selecting augmentation techniques is difficult, as there’s no fixed rule for their choice. This makes it necessary for researchers to choose data augmentation strategies carefully.
2.2. Generative Self-supervised Learning Methods
Inspired by the Mask language model (MLM) such as BERT (Qiu et al., 2021), generative methods take advantage of their powerful ability to process the sequence data when reconstructing corrupted data (i.e., sequences with masked items) to learn the embedding. In this scenario, $\hat{S}$ in the Eq. 5 usually is the set of data processed by the item masking method, and $L_{ssl}$ is the BERT-like loss function (e.g., Eq. 6). we categorize these methods based on their input types, with the majority of them designed to work with sequences, while only a limited subset is tailored for graph-based data.
2.2.1. Graph-based Generative Methods
It is difficult to directly conduct generative self-supervised learning using graph data. Thus, most existing work obtains sequences from graph data and then employs generative methods using the sequence data. For example, G-BERT (Shang et al., 2019) deals with the training data such that diagnosis and medical ontology are represented by a tree structure. The model utilizes the graph attention network (Veličković et al., 2017a) to learn the ontology embedding from diagnosis and medical tree-like graphs in order to make medication recommendations. Then, it feeds the ontology embeddings into a BERT-like model and pre-trains it with MLM-based tasks, which include the self-prediction task and the dual-prediction task. The former task is to predict the masked codes from the same type (i.e., medication or diagnosis) of ontology embeddings and the latter is to reconstruct the masked codes from different types of ontology embeddings.
PMGT (Liu et al., 2021d) focuses on multimodal recommendations, handling nodes with diverse information in a multimodal graph (e.g., item descriptions and images). They propose MCNSampling, a novel method to select neighboring nodes using importance scores and use BERT-Like patterns for node embedding learning. PT-GNN (Hao et al., 2021), instead of reconstructing masked nodes or items, addresses cold-start nodes with few neighbors using a meta-learning approach. It first pre-trains a transformer encoder on the embeddings of the first $K$ neighboring nodes to simulate cold-start scenarios. Then, a GCN model is pre-trained, incorporating the meta embedding in each convolution step. It finally maximizes cosine similarity between the predicted embedding and the ground-truth embedding generated by NCF (He et al., 2017) using all neighbor nodes.
2.2.2. Sequence-based Generative Methods
BERT4Rec (Sun et al., 2019) utilizes the BERT (Devlin et al., 2018) encoder for sequential recommendation. It uses BERT to model sequential user behaviors. BERT4Rec first applies item masking on the sample sequence, where the masked items are replaced with the special token ’[mask]’. Then it trains the model to predict the original item being masked based on the items nearby in the sequence. Its loss is defined as:
(6) $\mathcal{L}=\frac{1}{\left|\mathcal{V}_{u}^{m}\right|}\sum_{v_{m}\in\mathcal{V}_{u}^{m}}-\log P\left(v_{m}=v_{m}^{*}\mid\hat{\mathcal{S}_{u}}\right),$
where $\hat{\mathcal{S}_{u}}$ is obtained by applying item masking on the original user behavior sequence $\mathcal{S}_{u}$. $\mathcal{V}_{u}^{m}$ is the set of masked items. $v_{m}^{*}$ is the original item for the masked item $v_{m}$, and $P(v_{m}|\hat{\mathcal{S}_{u}})$ is the conditional probability calculated by the decoder of BERT. In the training phase, we can compute the item embedding through the encoder of BERT. There are several extensions of BERT4Rec for different tasks. For example, UNBERT (Zhang et al., 2021b) extends the BERT model for news recommendation. U-BERT (Qiu et al., 2021) aims to learn user embeddings of the target domains by transferring knowledge from content-rich source domains. UPRec (Xiao et al., 2021b) explores extending the BERT4Rec model to handle heterogeneous information such as user attributes and social networks.
Another line of research focuses on learning general item embeddings that are suitable for various recommendation tasks through generative methods. PeterRec (Yuan et al., 2020) is to conduct learning-to-learn idea (Bertinetto et al., 2016) on cross-domain recommendation task. Dissimilar to traditional cross-domain methods (Xie et al., 2022a; Wang et al., 2021c) which update the parameters of the whole model during the fine-tuning phase, it subtly inserts small neural network modules into the backbone of the original model and only trains the small neural networks to adapt to the downstream tasks. Like the application of BERT (Qiu et al., 2021) for learning the general words embedding, ShopperBERT (Shin et al., 2021) take advantage of the rich user behaviors to pre-train the BERT model based on nine auxiliary tasks for the general embedding. The experiment shows that it outperforms the model designed for one specific downstream recommendation task and indicates learning the general embedding is feasible.
In conclusion, generative methods take advantage of the successful paradigm of the MLM and the fitting ability of the transformer encoder resulting in learning the high performance and capacity of embedding. Although it is suitable for the sequence data naturally, researchers extend the methods to deal with graph embedding learning by transforming the graph data into the sequence data.
We summarize the data augmentation strategy, encoder, and recommendation task of all the self-supervised embedding learning methods mentioned in this section in Table 1.
Table 1. Summary of Self-supervised Embedding learning. ’GCNs’ denotes that the method indicates that the graph encoder could be instantiated as various GCN such as lightGCN (He et al., 2020), graph attention network.
Pretext Task Input Methods Data Augmentation Sample Encoder Recommendation Task
Contrastive Graph SGL (Wu et al., 2021b) Edge/Node Dropout, Random Walk GCN User Embedding/Node Embedding
HHGR (Zhang et al., 2021a) Node Dropout GCNs Group Recommendation
CCDR (Xie et al., 2022a) Subgraph Sampling GAT Cross-domain Recommendation
PCRec (Wang et al., 2021c) Subgraph Sampling GIN Cross-domain Recommendation
DCL (Liu et al., 2021c) Subgraph Sampling GCNs Next-item Recommendation
Sequence CL4SRec (Sun et al., 2019) Item Cropping/Reordering/Masking Transformer based Encoder Sequential Recommendation
CoSeRec (Liu et al., 2021a) Item Substitution/Insertion Transformer based Encoder Sequential Recommendation
ContraRec (Wang et al., 2022) Item Masking/Reordering Transformer based Encoder Sequential Recommendation
@@ 2.2.2. Sequence-based Generative Methods
Generative Graph G-BERT (Shang et al., 2019) Item Masking GCN/Transformer based Encoder Medical diagnosis Recommendation
PMGT (Liu et al., 2021d) Subgraph Sampling/Item Masking Transformer-based Graph Encoder Multimodal based Recommendation
PT-GNN (Hao et al., 2021) Subgraph Sampling GraphSAGE Cold-Start Problem
Sequence BERT4Rec (Sun et al., 2019) Item Masking Transformer based Encoder Sequential Recommendation
UNBERT (Zhang et al., 2021b) Word Masking Transformer based Encoder News Recommendation
U-BERT (Qiu et al., 2021) Word Masking Transformer based Encoder Cross-domain Recommendation
UPRec (Xiao et al., 2021b) Item Masking Transformer based Encoder Multimodal based Recommendation
2.3. Surveys and Tools
Surveys. Unlike this section only focuses on embedding learning for recommendation systems, Yu et al. (Yu et al., 2022) demonstrates self-supervised learning methods for the entire recommendation system. Furthermore, it provides a fine-grained taxonomy in SSL to introduce the recommendation methods. It also develops a toolkit for easy implementation of SSR methods.
Tools. SELFRec [^6] https://github.com/Coder-Yu/SELFRec is a framework for Self-supervised recommendation (SSR). It integrates multiple data augmentation techniques, popular datasets, and evaluation methods. Moreover, SELFRec is highly modularized and computationally efficient. It is also adaptable to the new SSR models.
2.4. Future direction
The choice of augmentation methods: Augmenting the data with a suitable and effective view is able to generate a strong self-supervised signal to optimize the model effectively. However, to search for workable augmentation techniques, scientists usually test many strategies randomly resulting in the waste of labor force and sub-optimal model performance. Several works (Xiao et al., 2020), (Kang et al., 2020b) has studied the theory of augmentation selection in contrastive learning but less in recommendation task. Thus, principles to guide the choice of augmentation methods in recommendation tasks are waiting for exploration.
On-device embedding for recommendation model: Recommendation models handle extensive user and item data, requiring substantial embedding capacity and memory. Traditional on-device deployment poses challenges due to device limitations. However, knowledge distillation has shown promise in shrinking model size, as demonstrated in works like (Wang et al., 2021e) and (Yu et al., 2021b). Integrating self-supervised learning methods could potentially compensate for performance loss in these compact models, an area with limited exploration.

### v3

@@ 2.3. Surveys and Future Directions
Survey. Abdi et al.(Abdi et al., 2018) focus on the field of context-aware recommendation based on matrix factorization, while Chen et al. (Chen et al., 2022a) analyze and conclude the theory of nonnegative matrix factorization with DNN. LibFFM [^2] LibFFM: https://github.com/ycjuan/libffm and xlearn333xlearn: https://github.com/aksnzhy/xlearn are used to achieve FFM (Juan et al., 2016). For other algorithms, corresponding toolkits can also be found. Such as DeepCTR [^4] DeepCTR: https://github.com/shenweichen/DeepCTR, which can achieve various CTR prediction models, including DeepFM (Guo et al., 2017).
@@
Future Direction: Synergistic Model Fusion for Complex Scenarios. An interesting avenue for future exploration lies in the synergy between different models. Instances like CAT-E and CAT-NT, as proposed by Zhao et al. (Zhao et al., 2017), creatively combine SVD, FM, and gradient boosting, showcasing the potential of model fusion. These approaches cater to highly specialized scenarios, often tied to the designs of MF models based on SVD and FM. Investigating and expanding upon such combinations could yield even more tailored solutions for intricate use cases. Except for the MF based on SVD & LFM, and FM, various kinds of CF development are also very meaningful for figuring out different specific scenarios in specific applications. For instance, DMF (Xue et al., 2017) uses explicit rating and non-preference implicit feedback to map the information of users and items to a low-dimensional space nonlinearly, which combines both the explicit and implicit information. EMF (Abdollahi and Nasraoui, 2016) combines the CF and graph weight, and XPL-CF (Almutairi et al., 2021) proposes to make the user-item relation clear.
3. Sequential Data Embedding
Sequential data, characterized by ordered interactions (e.g., user-item interaction sequences), is central to many recommender systems. Learning effective embeddings for such sequences requires capturing both temporal dependencies and higher-order patterns. Sequence embedding methods have evolved from RNN and CNN-based approaches to Transformer and state-space architectures, often enhanced with self-supervised learning (SSL) techniques. A detailed breakdown of these methods is provided in Table 3.
Table 2. Sequential Data Embedding
Sequential Data Embedding Sequential Modeling Techniques (§3.1) RNN/CNN-based Models: GRU4Rec (Hidasi et al., 2016), NARM (Li et al., 2017b), Caser (Tang and Wang, 2018), GLINT-RU  (Zhang et al., 2025)
Transformer-based Models: SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019)
Graph & State-Space Models: SR-GNN (Wu et al., 2019b), Mamba4rec (Liu et al., 2024a), STAR-Rec (Wang et al., 2025c), DiffuRec (Li et al., 2023a)
Contrastive Learning for Sequential Data Embeddings (§3.2) Key Models: CL4SRec (Xie et al., 2022b), CoSeRec (Liu et al., 2021a), ContraRec (Wang et al., 2022)
Generative Self-supervised Learning Methods (§3.3) Foundation Model: BERT4Rec (Sun et al., 2019; Devlin et al., 2018)
Model Extensions: UNBERT (Zhang et al., 2021b), U-BERT (Qiu et al., 2021), UPRec (Xiao et al., 2021b)
General Embedding Models: PeterRec (Yuan et al., 2020), ShopperBERT (Shin et al., 2021)
3.1. Sequential Modeling Techniques
Early sequence modeling in recommendation systems relied on recurrent or convolutional architectures. RNN-based models such as GRU4Rec (Hidasi et al., 2016) and NARM (Li et al., 2017b) encode sequences into hidden embeddings to capture temporal dependencies, while CNN-based models like Caser (Tang and Wang, 2018) use convolutional filters to capture local sequential patterns efficiently. Transformer-based models, including SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019), leverage self-attention to model long-term semantics, enhancing representation learning. Graph-based models such as SR-GNN (Wu et al., 2019b) construct graphs over sequence elements to capture higher-order transitions, and recent state-space or diffusion models (e.g., Mamba4rec (Liu et al., 2024a), DiffuRec (Li et al., 2023a)) efficiently model ultra-long or uncertain sequences. These architectures form the backbone of modern sequential recommendation, providing diverse ways to encode sequential dependencies and latent patterns in user-item interactions.
3.2. Contrastive Learning for Sequential Data Embeddings
@@
For sequence data, this section introduces four primary techniques to generate corrupted sequences to facilitate sequence augmentation. Item Cropping, inspired by image cropping in computer vision, is applied to contrastive learning in sequential recommendation. Item Masking, borrowed from language models, involves masking elements in a sequence to prevent overfitting and encourage learning of high-order structures among tokens. Similarly, Item Reordering is proposed to ensure predictions remain invariant to item order, training recommendation models on augmented data with varying item sequences. Item Substitution enhances recommendation models by maintaining the original sequence’s structure while substituting highly correlated yet possibly redundant items. Item Insertion, akin to substitution, involves adding related items before each original item in a subset of the sequence. Specific details are shown in Fig. 2.
Figure 2. Augmentation methods of Sequence. (a) Crop items for contrastive learning. (b) Mask item elements to prevent overfitting. (c) Augment data by reordering items, enhancing recommenders on varying sequences. (d) Substitute highly correlated yet possibly redundant items. (e) Add related items before the original item.
Following the general pipeline of CL, the CL for sequences first generates different views of each sequence by the augmentation methods mentioned above. Then train the recommendation model with contrastive loss (e.g., SimCLR (Chen et al., 2020a)). CL4SRec (Xie et al., 2022b) deals with the sequential recommendation task. It adopts three augmentation methods: item cropping, masking, and reordering. Given a minibatch of $N$ users, it generate 2NN augmented sequences $\{S_{u_{1}}^{a_{i}},S_{u_{1}}^{a_{j}},S_{u_{2}}^{a_{i}},S_{u_{2}}^{a_{j}},\ldots,S_{u_{N}}^{a_{i}},S_{u_{N}}^{a_{j}}\}$, where $S_{u_{n}}^{a_{i}}$ denotes the sequence interacted by user $u_{n}$ augmented by the augmentation method $a_{i}$ (e.g., item cropping). Then, for each user $u_{i}$, it regards $\left(S_{u_{i}}^{a_{i}},S_{u_{i}}^{a_{j}}\right)$ as the positive pair, and other 2(N1)2(N-1) elements as negative samples. Additionally, it utilizes the transformer encoder to compute the sequence embeddings. Eventually, the softmax cross-entropy loss is used to distinguish the positive pairs from the negative ones. Similar to CL4SRec, CoSeRec (Liu et al., 2021a) uses item substitution and insertion for data augmentation, while ContraRec (Wang et al., 2022) defines positive pairs from both the same sequence and different sequences with the same target item.
Contrastive methods utilize the CL loss with augmented data to enhance embeddings by capturing latent patterns. However, selecting augmentation techniques is difficult, as there’s no fixed rule for their choice. This makes it necessary for researchers to choose data augmentation strategies carefully. While contrastive methods effectively capture invariant patterns through augmentation, they face challenges in selecting appropriate transformation strategies. This limitation has led researchers to explore complementary approaches, particularly generative methods that learn through reconstruction rather than discrimination.
3.3. Generative Self-supervised Learning Methods
Generative methods in sequential recommendation leverage the principle of learning through reconstruction, inspired by the success of masked language modeling in NLP (Devlin et al., 2018). These approaches corrupt input sequences through masking or other perturbations, then train models to recover the original information, thereby learning robust sequential representations without explicit supervision. BERT4Rec (Sun et al., 2019) utilizes the BERT (Devlin et al., 2018) encoder for sequential recommendation. In particular, it uses BERT to model sequential user behaviors. BERT4Rec first applies item masking on the sample sequence, where the masked items are replaced with the special token ’[mask]’. Then it trains the model to predict the original item being masked based on the items nearby in the sequence. Its loss is defined as:
@@ 3.3. Generative Self-supervised Learning Methods
While these methods have demonstrated the effectiveness of basic SSL, a key challenge remains in understanding how SSL effectiveness varies across different recommendation contexts. Recent work has revealed significant performance variations depending on the specific recommendation scenario and data characteristics (Wei et al., 2023; Ren et al., 2024b). SSLRec (Ren et al., 2024b) provides a comprehensive framework enabling systematic comparison of SSL-enhanced models across five recommendation scenarios(e.g., multi-behavior and knowledge-enhanced recommendation). This unified evaluation reveals that different recommendation tasks require carefully tailored SSL strategies rather than one-size-fits-all approaches. MENTOR (Wei et al., 2023) demonstrates that multimodal recommendation benefits from simultaneous intra-modal and inter-modal contrastive objectives, where different information granularities require distinct contrastive strategies - visual features benefit from patch-level contrast while collaborative signals require user-item level contrast. The model utilizes the graph attention network (Veličković et al., 2017a) to learn the ontology embedding from diagnosis and medical tree-like graphs in order to make medication recommendations. The former task is to predict the masked codes from the same type (i.e., medication or diagnosis) of ontology embeddings, and the latter is to reconstruct the masked codes from different types of ontology embeddings. They propose MCNSampling, a novel method to select neighboring nodes using importance scores and use BERT-like patterns for node embedding learning. It first pre-trains a transformer encoder on the embeddings of the first $K$ neighboring nodes to simulate cold-start scenarios. Then, a GCN model is pre-trained, incorporating the meta embedding in each convolution step. It finally maximizes cosine similarity between the predicted embedding and the ground-truth embedding generated by NCF (He et al., 2017) using all neighbor nodes.
Although it is suitable for the sequence data naturally, researchers extend the methods to deal with graph embedding learning by transforming the graph data into sequence data. Thus, people should carefully consider the trade-off between the embedding quality and training cost.
3.4. Surveys and Future Directions
Surveys. While this section focuses on embedding learning for sequential data, several surveys provide comprehensive overviews of sequential recommendation methods (Fang et al., 2020; Dang et al., 2024; Pan et al., 2024; Boka et al., 2024). Pan et al.(Pan et al., 2024) offer an extensive survey on sequential recommendation, introducing a new taxonomy that categorizes various approaches, including ID-based, side information-enhanced, and generative models. Boka et al.(Boka et al., 2024) explore how sequential recommendation systems utilize interaction history to make more accurate and personalized recommendations. Future direction 1: The choice of augmentation methods. Augmenting the data with a suitable and effective view is able to generate a strong self-supervised signal to optimize the model effectively. However, to search for workable augmentation techniques, scientists usually test many strategies randomly, resulting in the waste of labor force and sub-optimal model performance. Several works (Xiao et al., 2020), (Kang et al., 2020b) have studied the theory of augmentation selection in contrastive learning, but less in the recommendation task. Thus, principles to guide the choice of augmentation methods in recommendation tasks are waiting for further exploration. Future direction 2: On-device embedding for recommendation model. Recommendation models handle extensive user and item data, requiring substantial embedding capacity and memory. Traditional on-device deployment poses challenges due to device limitations. However, knowledge distillation has shown promise in shrinking model size, as demonstrated in works like (Wang et al., 2021e) and (Yu et al., 2021). Integrating self-supervised learning methods could potentially compensate for performance loss in these compact models, an area with limited exploration. Thus, it is possible to learn a general-purpose embedding in a large-scale dataset that can serve different recommendation tasks directly or be fine-tuned with less computation cost. Some works (Zhang et al., 2020a), (Shin et al., 2021), (Yuan et al., 2020) have proposed some promising performance models, but they all take the BERT-like architecture. Therefore, it is a challenge to devise a new architecture for general-purpose embedding.
Table 3. Sequential Data Embedding
Sequential Data Embedding Sequential Modeling Techniques (§3.1) RNN/CNN-based Models: GRU4Rec (Hidasi et al., 2016), NARM (Li et al., 2017b), Caser (Tang and Wang, 2018), GLINT-RU  (Zhang et al., 2025)
Transformer-based Models: SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019)
Graph & State-Space Models: SR-GNN (Wu et al., 2019b), Mamba4rec (Liu et al., 2024a), STAR-Rec (Wang et al., 2025c), DiffuRec (Li et al., 2023a)
Contrastive Learning for Sequential Data Embeddings (§3.2) Key Models: CL4SRec (Xie et al., 2022b), CoSeRec (Liu et al., 2021a), ContraRec (Wang et al., 2022)
Generative Self-supervised Learning Methods (§3.3) Foundation Model: BERT4Rec (Sun et al., 2019; Devlin et al., 2018)
Model Extensions: UNBERT (Zhang et al., 2021b), U-BERT (Qiu et al., 2021), UPRec (Xiao et al., 2021b)
General Embedding Models: PeterRec (Yuan et al., 2020), ShopperBERT (Shin et al., 2021)